A disc rotates with a period of 0.50s. Its moment of inertia about its axis of rotation is 0.08 Kg m²

. A small mass is dropped onto the disc and rotates with it. The moment of 

inertia of the mass about the axis of rotation is 0.02 kg m²

. Determine the final period 

of the rotating disc and mass system.


1
Expert's answer
2021-01-19T07:11:39-0500

Let's apply the law of conservation of angular momentum:


"L_i=L_f,""I_{disc}\omega_{disc,i}+I_{mass}\omega_{mass,i}=(I_{disc}+I_{mass})\omega_{f}"

Since, initially small mass is at rest, we get:


"I_{disc}\omega_{disc,i}=(I_{disc}+I_{mass})\omega_{f}."

Taking into account that "\omega=\dfrac{2\pi}{T}," we get:


"I_{disc}\dfrac{2\pi}{T_{disc,i}}=(I_{disc}+I_{mass})\dfrac{2\pi}{T_{disc, f}},""T_{disc,f}=(\dfrac{I_{disc}}{T_{disc,i}(I_{disc}+I_{mass})})^{-1},""T_{disc,f}=(\dfrac{0.08\ kg\cdot m^2}{0.5\ s\cdot(0.08\ kg\cdot m^2+0.02\ kg\cdot m^2)})^{-1}=0.625\ s."

Answer:

"T_{disc,f}=0.625\ s."